OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? PLEASE HELP !!!!!!!!!!!

OpenStudy (anonymous):

@harsimran_hs4 yes i do know the slope intercept form is y=mx+c but how would you plug it in and get 2 equations?

OpenStudy (harsimran_hs4):

3 = m(-1) + c 11 = m(1) + c now can you solve for m and c?

OpenStudy (anonymous):

First you don't know the slope so you need to find the slop m=y2-y1/x2-x1

OpenStudy (anonymous):

okay i did that and got 8/2 .. correct? @begzat

OpenStudy (anonymous):

okay so next step y-y1=m*(x-x1) take one of the points and plug in to the equation

OpenStudy (phi):

okay i did that and got 8/2 .. correct? Yes, but 8/2 can be simplified to 4 that means the slope is 4. you now know y = 4x + b you need to find b

OpenStudy (anonymous):

how can i find b? @phi

OpenStudy (phi):

there are 2 ways (at least). one way is use begzat's formula another is replace y and x with numbers that you know are on the line. you know (-1, 3) is on the line. That means x= -1 and y=3 put those numbers in y= 4x + b 3= 4*-1+b can you solve for b?

OpenStudy (phi):

the first step is simplify 4* -1 is -4, so you have 3= -4 + b

OpenStudy (phi):

can you finish?

OpenStudy (anonymous):

sadly .. no i can't. @phi

OpenStudy (phi):

Ok, you might want to look at http://www.khanacademy.org/math/algebra/solving-linear-equations-and-inequalities/equations_beginner/v/simple-equations and the other videos. It is worth learning this or you will suffer in math class. Get some popcorn and watch them.

OpenStudy (anonymous):

but if he do with the formula that i told maybe will be much easy i think :)

OpenStudy (phi):

meanwhile 3= -4 + b you want the b by itself. To make that happen , we add +4 to both sides of the equation 3+4= -4+4+b the -4+4 =0 and the 3+4=7 7 = 0+b 0+b is just b we get 7= b or b=7 now you can write down the equation of the line

OpenStudy (anonymous):

Thank you so much and i will check out the videos. 4x - y = -7 correct? @phi

OpenStudy (phi):

we know m=4 and b=7, so y= 4x+7 if we move the y to the other side (add -y to both sides) y -y = 4x -y +7 or 0 = 4x -y +7 now move the 7 to the other side, (add -7 to both sides) -7 = 4x -y + 7 -7 or -7 =4x -y we can write that the other way around: 4x-y = -7 (notice you need to use algebra to swap the terms around)

OpenStudy (anonymous):

Thank you soo much! i'm studying for a really big test tomorrow lol thank you for the video link @phi